# LLM-ABBA: Understanding time series via symbolic approximation

### v3

@@ Abstract
Index Terms:
symbolic approximation, time series representation, quantization, time series regression, language models
@@ III-A1 Compression
Following the above, the whole polygonal chain can be recovered exactly from the first value $t_{0}$ and the tuple sequence $[p_{1},p_{2},\ldots,p_{N}]$ in the sense that the reconstruction error of this representation is with pinned start and end points and can be naturally modeled as a Brownian bridge. In terms of (1), a lower tol value is required to ensure an acceptable compression of time series with a great variety of features such as trends, seasonal and nonseasonal cycles, pulses and steps. As indicated in [13], the error bound between the reconstruction and original time series is upper bounded by $(n-N)\cdot\text{{tol}}^{2}$.
@@ III-B Error analysis reconstruction
$(\widehat{i}_{j},\widehat{t}_{i_{j}})=\bigg{(}\sum_{\ell=1}^{j}\widehat{\text{{len}}}_{\ell},t_{0}+\sum_{\ell=1}^{j}\widehat{\text{{inc}}}_{\ell}\bigg{)},\quad\text{for }j=0,\ldots,N,$ (3)
where $(\widehat{\text{{len}}}_{\ell},\widehat{\text{{inc}}}_{\ell})$ are the computed cluster centers, i.e., $\widehat{\text{{len}}}_{\ell}\in\mathcal{U}_{\text{{len}}}$ and $\widehat{\text{{inc}}}_{\ell}\in\mathcal{U}_{\text{{inc}}}$.
@@
$d^{\text{{inc}}}_{\ell}:=\widehat{\text{{inc}}}_{\ell}-\widetilde{\text{{inc}}}_{\ell},\quad d^{\text{{len}}}_{\ell}:=\widehat{\text{{len}}}_{\ell}-\widetilde{\text{{len}}}_{\ell}.$ (4)
@@ IV-A2 Hyperparameters of LLMs
Models Quant. Tokens Metric LoRA Optim. Epochs LR Batch Size
@@
Models Quant. Tokens Metric LoRA Optim. Epochs LR Batch Size
@@
Models Quant. Tokens Metric LoRA Optim. Epochs LR Batch Size
@@ IV-B Compression and Recovery
$1\times 10^{-2}$, $1\times 10^{-2}$ 3 846 $2.5\times 10^{-7}$ $1\times 10^{-2}$ $1.0$
$1\times 10^{-4}$, $1\times 10^{-4}$ 3 2,713 $4.2\times 10^{-8}$ $1.4\times 10^{-4}$ $1.0$
$1\times 10^{-6}$, $1\times 10^{-6}$ 3 2,789 $3.2\times 10^{-8}$ $1.3\times 10^{-4}$ 1.0
@@ IV-E Time Series Forecasting Tasks
Data Predictor Symbols Llama2-7B Mistreal-7B Informer [2] Time-LLM [6] TimeMixer [7]

### v4

@@ III-A1 Compression
Following the above, the whole polygonal chain can be recovered exactly from the first value $t_{0}$ and the tuple sequence $p_{1},p_{2},\ldots,p_{N}]$ in the sense that the reconstruction error of this representation is with pinned start and end points and can be naturally modeled as a Brownian bridge. In terms of (1), a lower tol value is required to ensure an acceptable compression of time series with a great variety of features such as trends, seasonal and nonseasonal cycles, pulses and steps. As indicated in [13], the error bound between the reconstruction and original time series is upper bounded by $(n-N)\cdot\text{{tol}}^{2}$.
@@ III-B Error analysis reconstruction
[…]
@@
$d^{\text{{inc}}}_{\ell}=\widehat{\text{{inc}}}_{\ell}-\widetilde{\text{{inc}}}_{\ell},\quad d^{\text{{len}}}_{\ell}=\widehat{\text{{len}}}_{\ell}-\widetilde{\text{{len}}}_{\ell}$. (4)
@@ IV-A2 Hyperparameters of LLMs
Models Quant. Tokens Metric LoRA Optim. Epochs
LR
Batch
Size
@@
Models Quant. Tokens Metric LoRA Optim. Epochs
LR
Batch
Size
@@
Models Quant. Tokens Metric LoRA Optim. Epochs
LR
Batch
Size
@@ IV-B Compression and Recovery
$\times 10^{-2}, 1×1$$\times 10^{-2}$ 3 846 $2.5\times 10^{-7}$ $1\times 10^{-2}$ $1.0$
$\times 10^{-4}, 1×1$$\times 10^{-4}$ 3 2,713 $4.2\times 10^{-8}$ $1.4\times 10^{-4}$ $1.0$
$\times 10^{-6}, 1×1$$\times 10^{-6}$ 3 2,789 $3.2\times 10^{-8}$ $1.3\times 10^{-4}$ 1.0
@@ IV-E Time Series Forecasting Tasks
Data Predictor Symbols Llama2-7B Mistreal-7B
Informer
[2]
Time-LLM
[6]
TimeMixer
[7]